we are stuck trying to solve this type of problem:
i don't understand how to simplify this expression:
10y to the seventh power over 6y squared

Answer by wuwei96815(245)

When you divide terms that have exponents, you subtract the exponents.

For example: 10y^7 / 6y^2 = 10/6y^5 = 5/3y^5 = 1 2/3y^5